# Direct measurement of atomic number density using Single Pass Absorption Spectroscopy (SPAS)

### v1

@@ Abstract
We demonstrate a direct measurement of the atomic number density of alkali atoms using single-pass absorption spectroscopy (SPAS). We developed our methodology based on modeling the absolute absorption spectra of warm rubidium (Rb) vapor and infer the atomic number density from SPAS measurements. The model combines the Lindblad formalism with a density matrix approach and incorporates experimentally measurable parameters such as laser beam power, laser beam diameter, and cell temperature. The framework explicitly incorporates optical pumping and transit-time broadening effects and shows excellent agreement ($99\%$) with experimental data using rubidium vapor cells across a wide range of temperature (293293343343 K), laser powers ($\sim 0.2~I_{sat}$$2~I_{sat}$), and cell lengths (22100100 mm). To ensure accurate quantification of absolute absorption measurements through the dilute atomic vapor, we measure and subtract the dark current of the photodetectors. This dark current is recorded in the absence of any light incident on the photodetector, to obtain accurate baseline corrections. This approach ensures an accurate determination of the atomic number density, even in the weak absorption regime. It provides a suitable alternative to the high-temperature, saturation-based method for baseline correction and enables the precise determination of the atomic number density in dilute atomic vapor for quantum technology applications in communication, sensing, and metrology using miniature atomic vapor cells. Furthermore, the methodology can be extended to determine the concentration of harmful gases and gaseous pollutants in urban, rural, as well as industrial environments.
Single-Pass Absorption Spectroscopy, Number density, Lindblad formalism, Density matrix approach, Rubidium vapor
@@ I Introduction
Precise determination of the number density of atoms is fundamental in atomic, molecular, and optical (AMO) physics, especially in applications like quantum sensing [1], frequency standards [2], laser cooling [3], and precision magnetometry [4]. In alkali atoms such as rubidium (Rb), the number density is strongly influenced by the transition strengths and absorption characteristics through changes in optical depth [5]. Naturally occurring Rb has two stable isotopes, 85Rb (72.2%) and 87Rb (27.8%) [6, 7], each exhibiting distinct hyperfine structures that are crucial for precision spectroscopy.
@@
Various experimental techniques have been employed to determine the atomic number density in warm alkali vapor cells. Fluorescence spectroscopy [8] is one such method, in which the population of the excited state can be inferred from the intensity of the emitted light. In contrast, Faraday rotation [9] correlates the polarization rotation of a probe beam with resonant atom density. Measurement of number density based on the Faraday rotation technique relies on spin polarization of the atomic sample and requires sophisticated magnetic field isolation using mu-metal shields as well as precision electronics for detection of spin polarization using precision balanced photodetectors. Thus, it is difficult to implement this method of number density measurement for out of laboratory field environments. The spin-exchange method [10] on the other hand infers the number density from collisional relaxation rates but is usually limited to specific regimes, e.g., high optical depth at substantially elevated temperatures, where direct absorption is impractical.
Despite the utility of these approaches, single pass absorption spectroscopy remains one of the most attractive, yet simple and one of the most widely used techniques globally to characterize atomic vapor density in near-room-temperature applications [11]. By monitoring the attenuation of a resonant probe beam as it traverses the atomic vapor and applying the Beer–Lambert law, one can indirectly extract the number density [12, 13]. Among various absorption-based techniques, Single-Pass Absorption Spectroscopy (SPAS) [14] stands out for its simplicity and effectiveness in probing the atomic transitions. However, while SPAS has been extensively utilized to investigate spectral line shapes and transition strengths, its explicit application for direct and quantitative determination of atomic number density has not been systematically reported.
@@
In this work, we introduce an approach to determine the total number density of a dilute atomic vapor using a a multi-level master equation approach. We validate our approach by modeling the single pass absorption spectra of rubidium $D_{2}$ transitions, utilizing the multi-level master equation framework integrating Doppler broadening within the framework. This method efficiently incorporates the intrinsic differences between the two stable isotopes, 85Rb and 87Rb, enabling accurate and computationally tractable modeling of their spectroscopic signatures. By treating both isotopes within a unified formalism, the model not only facilitates direct comparison with experimental data but also offers insights into the role of hyperfine structure, optical pumping, and relaxation mechanisms, such as transit-time broadening and finite laser linewidth. Importantly, our approach does not rely on conventional curve fitting of arbitrary parameters; instead, we incorporate all experimentally measured quantities, such as cell length, laser power, beam diameter, and temperature, into a theoretical model based on Lindblad formalism and extract the atomic number density as the only free parameter by matching the calculated absorption spectrum to the measured data. To the best of our knowledge, the work reported here represents the first measurement of the total number density of a dilute alkali vapor using single-pass absorption spectroscopy (SPAS) across a wide range of temperatures, laser beam powers, and different absorption lengths.
Furthermore, the given the model’s versatility to allow for a systematic exploration of absorption across a wide range of experimental parameters, including temperature, laser intensity, and interaction length, it makes our technique broadly applicable in the context of precision sensing and metrology. Although the present study focuses on the determination of the atomic number density in rubidium (Rb) atoms, our approach is readily extendable to other atomic and molecular species as well. Thus, our technique holds promise for practical applications, for instance, in detecting trace gases or harmful pollutants in field environments, where an analysis of the single pass laser absorption profile can yield quantitative information on isotopic composition and number density.
@@
The manuscript is organized as follows. In section II, we outline the theoretical framework for modeling the absorption spectra, detailing the development of a multi-level master equation approach and integrating Doppler broadening into absorption coefficient calculations. In section III we describe the experimental methodology, including the layout of the optical setup and the techniques used for signal acquisition. In the following section section IV, we present a comparative analysis between the simulated and experimentally observed spectra, emphasizing the influence of critical parameters such as temperature and laser power in the number density estimation. Finally, in the last section V, we conclude our studies by summarizing key findings and discussing potential avenues to extend this approach to other atomic and molecular systems for use in sensing and metrology applications.
@@ II Theoretical background
In the presence of light–matter interaction and dissipative processes such as dephasing and spontaneous emission, the state vector alone is no longer sufficient to capture the complete dynamics of the system [15]. Instead, the density matrix denoted by $\rho$ is used because it can describe both pure and mixed states, including decoherence, spontaneous emission, and other dissipative effects [16, 17, 18]. For an $N$-level quantum system, the density matrix is an $N\times N$ Hermitian matrix. In this matrix, the diagonal elements $\rho_{ii}$ indicate the probability (or population) of finding the system in the $i$-th state. In contrast, the off-diagonal elements $\rho_{ij}$ represent the quantum coherence between the states $|i\rangle$ and $|j\rangle$. This matrix framework is essential in scenarios where the quantum system interacts with external fields or environments, leading to a loss of coherence.
In this work, we use the density matrix formalism, where the atom is considered as a quantum mechanical system and the electric field is assumed to be a classical field. We construct the Hamiltonian of the Rb $D_{2}$ transition by including relevant atomic energy levels and their interaction with the light field. We solve the dynamics of the system using the master equation. The results are an accurate depiction of the coherence terms, including both absorption and dispersion. The approach is semiclassical, as we are considering the electric field to be classical, remains computationally less intensive than a complete quantum treatment of both the atom and the field, as it avoids the complexities associated with a fully quantized electric field, and still provides a reliable and quantitative understanding of the dynamics of the system.
@@
To derive such a Hamiltonian, we use the semi-classical approach [19, 20] in which the atom is treated quantum mechanically while the radiation field is taken as a classical oscillating field. The interaction Hamiltonian has the general form $H_{I}(t)=-\textbf{{d}}\cdot\textbf{{E}(t)}$, where d is the electric dipole operator and E(t) is the electric field. Expand the dipole operator in the atomic basis and introduce the Rabi frequency $\Omega_{ij}=\frac{E_{0}}{\hbar}D_{ij}$ [21], which quantifies the strength of the coupling between the ground $|i\rangle$ and excited states $|j\rangle$ (by definition $\Omega_{ij}=\Omega_{ji}^{*}$ and $\Omega_{ii}=0$), where $E_{0}$ is the field amplitude and $D_{ij}$ is the dipole matrix element [22]. By applying the rotating-wave approximation (RWA) to keep only the absorption $|i\rangle\to|j\rangle$ and emission $|j\rangle\to|i\rangle$ terms, and then transforming into the rotating frame to get the time-independent Hamiltonian of the four-level system,
@@
To simulate the absorption spectroscopy of the rubidium $D_{2}$ transition, we construct a model using two separate four-level atomic systems, for each isotope: $^{85}\mathrm{Rb}$ and $^{87}\mathrm{Rb}$. Each system includes a ground state and three excited states. For a given isotope, the model distinguishes between transitions (dipole allowed) originating from the lower ($|0_{a}\rangle$) and upper ($|0_{b}\rangle$) hyperfine ground states, which couple to the common excited states $|2\rangle$ and $|3\rangle$, and to the exclusive excited states $|1\rangle$ and $|4\rangle$, respectively, as shown in Figure 1(b). This framework enables accurate modeling of hyperfine population dynamics, optical pumping effects, and decay pathways, incorporating the experimentally determined branching ratios between states.
@@ II.2 Electrical susceptibility and absorption coefficient
In our theoretical model, we explicitly consider the hyperfine structure of the rubidium $D_{2}$ line ($5^{2}S_{1/2}\leftrightarrow 5^{2}P_{3/2}$), constructing two distinct four-level systems for the two naturally occurring isotopes, 87Rb and 85Rb. As shown in Figure 1(b) for 87Rb, the hyperfine levels of the ground state are $|0_{a}\rangle=|F=1\rangle$ and $|0_{b}\rangle=|F=2\rangle$, while the hyperfine excited state ($5^{2}P_{3/2}$) consists of $\lvert 1\rangle$ to $\lvert 4\rangle$, corresponding to $\lvert F^{\prime}=0\rangle$ to $\lvert F^{\prime}=3\rangle$ respectively. For 85Rb, the ground state levels are $|0_{a}\rangle=\lvert F=2\rangle$ and $\lvert 0_{b}\rangle=\lvert F=3\rangle$ , with excited state levels $\lvert 1\rangle$ to $\lvert 4\ranglecorresponding$ to $\lvert F^{\prime}=1\rangle$ to $\lvert F^{\prime}=4\rangle$ respectively.
@@
where $C_{f}{i\rightarrow f}$ is the transition strength from the initial state $\lvert i\rangle$ to a particular final state $\lvert f\rangle$. The branching ratios corresponding to all allowed transitions of the $D_{2}$ line of both 85Rb and 87Rb are presented in the Table 1.
@@ II.3 Transit time and Doppler broadening
The linewidth of an absorption spectrum in an atomic vapor cell is influenced by several homogeneous and inhomogeneous broadening mechanisms, among which transit-time relaxation and Doppler broadening play significant roles. The transit-time relaxation arises from the interaction time between moving atoms and the incident laser beam. As atoms with velocities from the Maxwell-Boltzmann distribution traverse the probe region, they are effectively reset upon entering and exiting. When an atom leaves the interaction volume, it is statistically replaced by a new, unpumped atom from the surrounding ensemble. From the system’s perspective, this process acts as an additional relaxation pathway that contributes to linewidth broadening. Unlike spontaneous emission, which only occurs from excited states to ground states, transit-time relaxation introduces a dephasing mechanism that affects both ground and excited states uniformly. It destroys the phase coherence of the atomic ensemble, thereby broadening the spectral line. The transit relaxation rate, denoted as $\gamma_{t}$ expressed as,
@@
To obtain the Doppler broadened spectra, the absorption coefficient from Equation 7 is integrated over all atomic velocities,
@@
Ideally, the simulation of the Doppler broadened absorption coefficient (in Equation 9) should span a velocity range that captures nearly all atoms in the vapor. However, to balance efficiency and computation time, we consider the velocity range from $4\sigma_{v}$ to $4\sigma_{v}$, , which includes approximately $99.99\%$ of the atomic population. Here, $\sigma_{v}=\sqrt{\frac{k_{B}T}{m}}$ denotes the root-mean-square (rms) velocity of the atoms along the laser propagation direction, where $k_{B}$ is the Boltzmann constant, $T$ is the temperature of the atomic vapor, and mm is the atomic mass.

### v2

# Direct determination of atomic number density in MEMS vapor cells via single-pass absorption spectroscopy (SPAS)

@@ Abstract
Micro-electro-mechanical systems (MEMS)-based chip-scale alkali vapor cells are the essential components in emerging quantum technologies, including compact atomic clocks, chip-scale magnetometers, and miniature quantum opto-electronic systems. The sensitivity of MEMS-based atomic quantum technology devices depends on the atomic number density. Thus, it is important to have an accurate estimate of the atomic number density in chip-scale alkali vapor cells to optimize light-matter interactions and design efficient quantum sensing systems. Here, we present a direct and quantitatively validated method for determining the rubidium (Rb) number density in warm alkali vapor using single-pass absorption spectroscopy (SPAS). The absolute transmission spectra are measured and modeled using the 780.24 nm as well as the 420.29 nm transition in a Rb-filled MEMS vapor cell. The atomic number density measurements and the model were also validated using a commercial vapor cell of length 100 mm. The theoretical model employs a density-matrix formalism within the Lindblad framework and incorporates directly measurable experimental parameters, such as laser beam power, diameter, and cell temperature. The model explicitly accounts for optical pumping, Doppler broadening, and transit-time broadening effects and exhibits quantitative agreement ($99\%$) with experimental spectra over a broad range of temperatures (293–343 K), laser intensities ($\sim 0.2\,I_{\mathrm{sat}}$ to $\sim 2\,I_{\mathrm{sat}}$), and cell lengths (2–100 mm). The extracted densities from the MEMS cell closely follow the empirical vapor-pressure model by Alcock et al. The demonstrated methodology provides a practical, well-controlled method for determining the atomic number density in alkali vapor cells relevant to the characterization and development of compact alkali-vapor-based devices for quantum sensing and metrology.
Single-Pass Absorption Spectroscopy, number density, Density matrix, MEMS
@@ I Introduction
With the growing demand for portable quantum devices for use in precision timing, sensitive magnetometry [1], electric-field sensing [2, 3, 4] , and entangled photon-pair generation [5, 6], etc., micro-fabricated chip-scale vapor cells based on micro-electro-mechanical systems (MEMS) technology [7] have become an indispensable tool in the domain of field-deployable atomic quantum technology devices for use in quantum sensing and quantum communication applications. These miniaturized vapor cells enable wide applications in quantum technologies, including compact atomic clocks [8], high-precision magnetometers [9], frequency-metrology platforms [10], and integrated photonic–atomic sensors [11]. Their compact form factor, low power consumption, greater reliability, and low cost for scalable fabrication make them particularly attractive for next-generation quantum applications. Most quantum devices rely on the well-defined optical and spectroscopic properties of alkali vapors, enabling applications ranging from inertial navigation [12] and timing synchronization [13] to biomedical applications through ultra-low magnetic field detection [14].
In chip-scale quantum devices, particularly MEMS-based vapor cells, the intrinsically short optical path length makes the optical depth extremely sensitive to the atomic number density $N$. Since the stability and sensitivity of many atomic and quantum technologies scale as $1/\sqrt{\mathcal{N}}$, where $\mathcal{N}$ is the effective number of atoms/ions participating in the light–matter interaction [15, 16], precise knowledge of the ensemble density is essential for accurate device design for optimized performance and quantitative modeling. Reliable determination of number density is therefore critical for optimizing the performance of miniature atomic clocks, optically pumped magnetometers, photon-pair sources, nonlinear optical interfaces, and quantum memory systems, while avoiding density-dependent systematic shifts and unwanted broadening effects.
In alkali atoms, rubidium (Rb) is one of the most widely used platforms for MEMS-based vapor-cell applications due to its well-defined spectroscopic structure, strong optical transitions, and accessible vapor pressures near room temperature [17]. Naturally occurring Rb has two stable isotopes, 85Rb (72.2%) [18] and 87Rb (27.8%) [19], each exhibiting distinct multi-level hyperfine structures that must be accurately modeled to extract quantitative information of the atomic number density. Estimating the number density of warm alkali vapors typically relies on empirical vapor-pressure relations [20]. The standard expression for the temperature-dependent vapor pressure $P_{\mathrm{vap}}(T)$ was first proposed by Killian [21] and later refined by Alcock, Itkin, and Horrigan [20]. For rubidium, the vapor pressure $P_{\mathrm{vap}}(T)$ (in torr) as a function of temperature $T$ (in Kelvin) is given by
$\log_{10}P_{\mathrm{vap}}(T)=2.881+a-\frac{b}{T},$ (1)
where the constants $a$ and $b$ correspond to different physical phases: $a=4.857,b=4215~\text{(solid phase)},a=4.312,b=4040~\text{(liquid phase)}.$ Here, $P_{\mathrm{vap}}$ is expressed in torr; when expressed in atmospheres, the additive constant 2.881 is omitted. The corresponding atomic number density $N(T)$ is determined from the ideal-gas relation,
$N(T)=\frac{P_{\mathrm{vap}}(T)}{k_{\mathrm{B}}T},$ (2)
where $k_{\mathrm{B}}$ is the Boltzmann constant and $P_{\mathrm{vap}}$ is converted to Pascals using $1~\text{torr}=133.322~\text{Pa}$. This empirical formulation remains the standard benchmark for validating experimentally inferred alkali-vapor densities.
@@
Various experimental techniques have been employed to determine the atomic number density in warm alkali vapor cells, including fluorescence spectroscopy [22], Faraday rotation [23], and spin-exchange measurements [24]. In fluorescence spectroscopy, the population of the excited state can be inferred from the intensity of the emitted light. In contrast, Faraday rotation correlates the polarization rotation of a probe beam with resonant atom density. Measurement of the number density based on the Faraday rotation technique relies on the spin polarization of the atomic sample. It requires sophisticated magnetic-field isolation using mu-metal shields, as well as precision electronics for detecting spin polarization with balanced photodetectors. Thus, implementing this method of number density measurement in field environments outside the laboratory is challenging. The spin-exchange method infers the number density from collisional relaxation rates. Still, it is usually limited to specific regimes, e.g., high optical depth at substantially elevated temperatures, where direct absorption is impractical. Despite the utility of these approaches, single-pass absorption spectroscopy (SPAS) [25] remains one of the most attractive, simple, and widely accessible methods for characterizing warm vapor systems, especially near room temperature [26]. However, while SPAS has been extensively used to investigate spectral lineshapes [27] and transition strengths [26], it has rarely been used for the explicit, quantitative determination of absolute number density in a MEMS-based vapor cell. This limitation arises from the need for accurate baseline calibration and the complexities introduced by optical pumping, multi-level dynamics, and Doppler broadening in chip-scale atomic vapor cells.
@@
In this work, we present a systematic, quantitatively validated framework for determining the atomic number density of a dilute atomic vapor using a multi-level master equation model. Our theoretical approach efficiently incorporates the intrinsic differences between the two stable isotopes, 85Rb and 87Rb, enabling accurate and computationally tractable modeling of their spectroscopic signatures. By treating both isotopes within a unified formalism, the model not only facilitates direct comparison with experimental data but also offers insights into the role of hyperfine structure, optical pumping, and relaxation mechanisms, such as transit-time broadening and finite laser linewidth. Importantly, our approach does not rely on conventional curve fitting of arbitrary parameters; instead, we incorporate all experimentally measured quantities, such as cell length, laser power, beam diameter, and temperature, into a theoretical model based on Lindblad formalism and extract the atomic number density as the only free parameter by matching the calculated absorption spectrum to the measured data. To the best of our knowledge, the work reported here represents the first measurement of the total number density of a dilute alkali vapor using single-pass absorption spectroscopy across a wide range of temperatures, laser beam powers, and different absorption lengths.
We demonstrate this method across markedly different optical path lengths, using both a MEMS vapor cell (2 mm, designed and fabricated in the LEOS (Laboratory for Electro-Optics Systems), ISRO (Indian Space Research Organization)[28]) and a conventional 100 mm (from Triad Technology Inc.) vapor cell, and at two widely separated wavelengths corresponding to the $5S_{1/2}\rightarrow 5P_{3/2}$ (780.24 nm) and $5S_{1/2}\rightarrow 6P_{3/2}$ (420.29 nm) transitions. This broad validation highlights the generality of the technique and its relevance for precision quantum-optics experiments, spectroscopy in weak-absorption regimes, and the characterization of chip-scale alkali-vapor devices. Although our study focuses on Rb, the approach is readily extendable to other atomic and molecular species, making it a versatile tool for laboratory and field-deployable sensing applications.
@@
The manuscript is organized as follows. In section II, we outline the theoretical framework for modeling the absorption spectra, detailing the development of a multi-level master equation approach and integrating Doppler broadening into absorption coefficient calculations. In section III we describe the experimental methodology, including the layout of the optical setup and the techniques used for signal acquisition. In the following section section IV, we present a comparative analysis of the simulated and experimentally observed spectra, emphasizing the influence of critical parameters, such as temperature and laser power, on the number density estimation. Finally, in section V, we summarize key findings and discuss potential avenues to extend this approach to other atomic and molecular systems for sensing and metrology applications.
@@ II Theoretical background
In the presence of light–matter interaction and dissipative processes such as dephasing and spontaneous emission, the state vector alone is no longer sufficient to capture the complete dynamics of the system [29]. Instead, the density matrix denoted by $\rho$ is used because it can describe both pure and mixed states, including decoherence, spontaneous emission, and other dissipative effects [30, 31, 32]. For an $N$-level quantum system, the density matrix is an $N\times N$ Hermitian matrix. In this matrix, the diagonal elements $\rho_{ii}$ indicate the probability (or population) of finding the system in the $i$-th state, and the off-diagonal elements $\rho_{ij}$ represent the quantum coherence between the states $|i\rangle$ and $|j\rangle$. This matrix framework is essential in scenarios where a quantum system interacts with external fields or environments, leading to coherence loss.
Figure 1: Energy-level diagram of rubidium (87Rb) showing the 420 nm excitation scheme, relevant intermediate states, and decay pathways, highlighting the 780 nm and 420 nm transitions with their corresponding branching ratios [33] used to determine the atomic number density of the Rb atoms in the MEMS vapor cell.
In this work, we use the density-matrix formalism, in which the atom is treated quantum-mechanically, and the electric field is treated classically. We construct the Hamiltonian of a four-level system for the transition of Rb atom from ground state $5S_{1/2}$ to the excited state $5P_{3/2},6P_{3/2}$ by including the relevant atomic energy levels (as shown in Figure 1) and their interactions with the light field. We solve the steady state of the system using the master equation. The results accurately depict the coherence terms, including both absorption and dispersion. To simulate the absorption spectroscopy of these Rb transitions ($5S_{1/2}\to 5P_{3/2},6P_{3/2}$), we construct a theoretical model using two separate four-level atomic systems, for each isotope: $^{85}\mathrm{Rb}$ and $^{87}\mathrm{Rb}$. Each system includes a ground state and three excited states. For a given isotope, the model distinguishes between transitions (dipole allowed) originating from the lower ($|0_{a}\rangle$) and upper ($|0_{b}\rangle$) hyperfine ground states, which couple to the common excited states $|2\rangle$ and $|3\rangle$, and to the exclusive excited states $|1\rangle$ and $|4\rangle$, respectively, as shown in Figure 2(b). This framework enables accurate modeling of optical pumping effects and decay pathways, incorporating the branching ratios between states.
@@
To derive such a Hamiltonian [34, 35]. The interaction Hamiltonian has the general form $H_{I}(t)=-\textbf{{d}}\cdot\textbf{{E}(t)}$, where d is the electric dipole operator and E(t) is the electric field. Expand the dipole operator in the atomic basis and introduce the Rabi frequency $\Omega_{ij}=\frac{E_{0}}{\hbar}D_{ij}$ [36], which quantifies the strength of the coupling between the ground $|i\rangle$ and excited states $|j\rangle$ (by definition $\Omega_{ij}=\Omega_{ji}^{*}$ and $\Omega_{ii}=0$), where $E_{0}$ is the field amplitude and $D_{ij}$ is the dipole matrix element [37]. By applying the rotating-wave approximation (RWA) to keep only the absorption $|i\rangle\to|j\rangle$ and emission $|j\rangle\to|i\rangle$ terms, and then transforming into the rotating frame to get the time-independent Hamiltonian of the four-level system,
@@ II.2 Electrical susceptibility and absorption coefficient
In our theoretical model, we explicitly consider the hyperfine structure of the rubidium atom at 780 nm and 420 nm ($5^{2}S_{1/2}\leftrightarrow 5^{2}P_{3/2},6^{2}P_{3/2}$ ) transitions, constructing two distinct four-level systems for the two naturally occurring isotopes, 87Rb and 85Rb. As shown in Figure 2(b) for 87Rb, the hyperfine levels of the ground state are $|0_{a}\rangle=|F=1\rangle$ and $|0_{b}\rangle=|F=2\rangle$, while the hyperfine excited state ($5^{2}P_{3/2},6^{2}P_{3/2}$) consists of $\lvert 1\rangle$ to $\lvert 4\rangle$, corresponding to $\lvert F^{\prime}=0\rangle$ to $\lvert F^{\prime}=3\rangle$ respectively. For 85Rb, the ground state levels are $|0_{a}\rangle=\lvert F=2\rangle$ and $\lvert 0_{b}\rangle=\lvert F=3\rangle$ , with excited state levels $\lvert 1\rangle$ to $\lvert 4\ranglecorresponding$ to $\lvert F^{\prime}=1\rangle$ to $\lvert F^{\prime}=4\rangle$ respectively.
@@
where $C_{f}{i\rightarrow f}$ is the transition strength from the initial state $\lvert i\rangle$ to a particular final state $\lvert f\rangle$. Since the total electronic angular momentum $J$, nuclear spin $I$, and orbital angular momentum $L$ are identical for the $5P_{3/2}$ and $6P_{3/2}$ states, the hyperfine structure, Clebsch–Gordan coefficients, and branching ratios for individual hyperfine transitions remain unchanged relative to the $D_{2}$ line. Despite these similarities, the overall absorption strength of the $5S_{1/2}\rightarrow 6P_{3/2}$ transition is significantly weaker than that of the $D_{2}$ transition. Unlike the $D_{2}$ line, where spontaneous decay from the excited state occurs exclusively to the ground state, the $6P_{3/2}$ state exhibits multiple radiative decay channels, as shown in Figure 1. Approximately $23\%$ of the total spontaneous decay from $6P_{3/2}$ returns directly to the ground state $5S_{1/2}$ [33], with the remaining population decaying via intermediate states. This reduced branching ratio to the ground state and a smaller dipole matrix element led to substantially weaker absorption on the $420~\mathrm{nm}$ transition. The branching ratios corresponding to all allowed transitions of both 85Rb and 87Rb are presented in Table 1.
@@ II.3 Transit time and Doppler broadening
The linewidth of an absorption spectrum in an atomic vapor cell is influenced by several homogeneous and inhomogeneous broadening mechanisms, among which transit-time relaxation and Doppler broadening play significant roles. The transit-time relaxation arises from the interaction time between moving atoms and the incident laser beam. As atoms with velocities from the Maxwell-Boltzmann distribution traverse the probe region, they are effectively reset upon entering and exiting. When an atom leaves the interaction volume, it is statistically replaced by a new, unpumped atom from the surrounding ensemble. From the system’s perspective, this process acts as an additional relaxation pathway that contributes to linewidth broadening. Unlike spontaneous emission, which occurs only from excited states to the ground state, transit-time relaxation introduces a dephasing mechanism that affects both ground and excited states uniformly. It destroys the phase coherence of the atomic ensemble, thereby broadening the spectral line. The transit relaxation rate, denoted as $\gamma_{t}$ expressed as,
@@
To obtain the Doppler-broadened spectra, the absorption coefficient from Equation 9 is integrated over all atomic velocities,
@@
Ideally, the simulation of the Doppler-broadened absorption coefficient (in Equation 11) should span a velocity range that captures nearly all atoms in the vapor. However, to balance efficiency and computation time, we consider the velocity range from $4\sigma_{v}$ to $4\sigma_{v}$, which includes approximately $99.99\%$ of the atomic population. Here, $\sigma_{v}=\sqrt{\frac{k_{B}T}{m}}$ denotes the root-mean-square (rms) velocity of the atoms along the laser propagation direction, where $k_{B}$ is the Boltzmann constant, $T$ is the temperature of the atomic vapor, and mm is the atomic mass.